Write a pair of equivalent equations which can be solved by adding or subtracting
1) 2x+3y=10
3x+4y=-2

2) 7x-6=10
2x+5y=23

Answer by ikleyn(52835)

.
Write a pair of equivalent equations which can be solved by adding or subtracting
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
1)        

2x + 3y = 10,   (1)
3x + 4y = -2.   (2) 

Multiply the equation (1) by 3 (both sides) and the equation (2) by 2. You will get an equivalent system

6x + 9y = 30,   (1')
6x + 8y = -4    (2')

The last system can be solved by distraction.


2)        

7x - 6  = 10,
2x + 5y = 23.
Are you sure you correctly wrote this system?